4. Expand the following functions in a Maclaurin series: (2) 1+x (b) (1 + x)? (e) (1 + x)12 (d) In(1 - x) (e) er

f) a^x
Answer #1

I 1+x The Maclaurin Series is, FC = : f0] Hflosz+f%b2%. fW f) for ito !, flessitors f(x) .fo = ca 2 f = 33 de o filho - S9 . 2 CHxby Fax) = 1-2x + 30% - 4x3+ -- for = (1+3) Ya . . f(0) = fou = 1 (1+x) : fb) = { f(x) = 1 Clts = C.1) (142)! -] (f = 1 (1-x) = . - 1 (-1261) - Le Cl-x)2 = (1-x)2 f(o) F(x) = 1 2 = (1-x B 23 Noce, f (0) = 2. +(x) = lut + 1.X + 2.3° t -